Solve system by elimination
1/2x+y=22
2x+4y=11

Answer by elima(1433)

1/2x+y=22
2x+4y=11
multiply -4 to the first equation;
(-4)1/2x+y=22
-2x-4y=-88
now add the first equation and second together;
-2x-4y=-88
2x+4y=11
-----------
0=-77
no solution
:)